OpenStudy (anonymous):

y(exponent 5)*y(exponent 7) idk how to type exponents. my bad.

OpenStudy (anonymous):

y^5*y^7 = y^(5+7) = y12 , regarding the way you may edit method , it's either trough using this notation ^ or through the Equation editor you find at the bottom of your editing box

OpenStudy (anonymous):

right... so you pretty much just add up the exponents?

OpenStudy (anonymous):

\[ y^{5} * y^{7} = y^{5+7} = y^{12}\]

OpenStudy (anonymous):

perfect... thanks

OpenStudy (anonymous):

yw